OpenStudy (anonymous):

if a^2+b^2=7ab then log(a+b)/3 is equal to

OpenStudy (anonymous):

a^2+b^2+2ab = 9ab (a+b)^2=9ab taking log 2log (a+b )= log 9 + log a +log b 2log(a+b)-2log3 = log a + log b 2log(a+b)/3= log a + log b log (a+b)/3= log(ab)/2

OpenStudy (dumbcow):

oh i didnt realize the 3 was inside the log :{

OpenStudy (anonymous):

thankz all